# Amplifying hybrid entangled states and superpositions of coherent states

### v1

@@ III Amplification of Hybrid State
For the $\hat{a}\hat{a}^{\dagger}$ scheme, solving the equation gives the gain as
$g=\frac{\alpha^{2}-1+\sqrt{\alpha^{4}+6\alpha^{2}+1}}{2\alpha^{2}},$ (20)
while for the $(\hat{a}^{\dagger})^{2}$ scheme, solving the equation gives the gain as
$g=\frac{1+\sqrt{1+\frac{8}{\alpha^{2}}}}{2}.$ (21)
The gain graphs of the two schemes are shown in Fig 1. (b). One can see that the $(\hat{a}^{\dagger})^{2}$ scheme yields higher gain than the $\hat{a}\hat{a}^{\dagger}$ scheme, converse to the fidelity graphs. The behavior of optimal gain of hybrid state amplification is also identical to that of coherent state amplification for the same reason as fidelity.
We further note that the gain can be defined by the ratio of the expectation values of quadrature operators with an arbitrary phase $\phi$ Zavatta11 ; Park16 . This gain value, however, can be somewhat inadequate for hybrid states because the expectation value of quadrature value with any phase $\phi$ is zero for hybrid states, as one can confirm with the following equation.
@@
$g=\frac{(N^{f(\hat{a},{\hat{a}^{\dagger}})})^{2}\bra{\alpha}f(\hat{a},{\hat{a}^{\dagger}})^{\dagger}(\hat{a}e^{i\lambda}+\hat{a}^{\dagger}e^{-i\lambda})f(\hat{a},{\hat{a}^{\dagger}})\ket{\alpha}}{\bra{\alpha}(\hat{a}e^{i\lambda}+\hat{a}^{\dagger}e^{-i\lambda})\ket{\alpha}}$
Let us writhe the normal ordered form of $f(\hat{a},{\hat{a}^{\dagger}})^{\dagger}\hat{a}f(\hat{a},{\hat{a}^{\dagger}})$ as $\sum_{i}C_{i}\,\hat{a}^{\dagger}{}^{i}\hat{a}^{\dagger}{}^{i}\hat{a}=g(\hat{a},\hat{a}^{\dagger})\hat{a}$, then the gain is written as
$g=(N^{f(\hat{a},{\hat{a}^{\dagger}})})^{2}\bra{\alpha}g(\hat{a},\hat{a}^{\dagger})\ket{\alpha}$ (23)
Likewise, the gain of hybrid states defined by the ratio is
$g_{hyb}=(N_{hyb}^{f(\hat{a},{\hat{a}^{\dagger}})})^{2}\bra{\alpha}g(\hat{a},\hat{a}^{\dagger})\ket{\alpha}$
@@
We can go further with this gain to define equivalent input noise (EIN), a measure of noiseless amplification. However, since the expectation value of quadrature is zero, it lacks physical meaning, thus we will not stick to it.
@@
Observing the results so far, $\hat{a}\hat{a}^{\dagger}$ scheme has an advantage over higher fidelity while $(\hat{a}^{\dagger})^{2}$ has an advantage over higher gain, for all amplitude $\alpha$. This property gives us a guide on which amplification scheme to use. However, their advantages over different aspects make it hard to compare their overall performance. Qualitatively, we know that $\hat{a}\hat{a}^{\dagger}$ scheme would be better for small amplitude since coherent states with small amplitude are susceptible to the change of their photon number. Meanwhile, we can guess that the two schemes will not differ so much for large amplitude because coherent states with sufficiently large amplitude are robust against several photon addition or subtraction. Lastly, as a consequence, we can deduce that the $(\hat{a}^{\dagger})^{2}$ scheme would be appropriate for the intermediate amplitude region.
@@
Previous research Park16 employed two measures, EIN for coherent states and quantum Fisher information for $2$-dimensional SCSs to quantify overall performance which reflects aforementioned qualitative arguments. EIN is a measure of undesired noise generated in the amplification process, thus it is appropriate to measure overall performance. However, we have concluded that EIN lacks physical meaning in our case since the expectation values of quadrature operators are always zero. As an alternative, we take quantum Fisher information for phase estimation as a figure of merit of amplified states, which is defined by
$\mathcal{F}[\rho,\hat{H}]=2\sum_{k,l}\frac{(\lambda_{k}-\lambda_{l})^{2}}{(\lambda_{k}+\lambda_{l})}|\bra{k}\hat{H}\ket{l}|^{2},$ (24)
@@
where $\rho$ is a density matrix, |$\ket{k}$ is a $k$-th eigenvector and $\lambda_{k}$ is a $k$-th eigenvalue of the density matrix $\rho$, $\hat{H}$ is the Hamiltonian operator of the harmonic oscillator. Quantum Fischer information is in direct relationship with the performance of phase estimation, known as the quantum Cramér-Rao bound.
$\mathrm{var}(\hat{\theta})\geq\frac{1}{\mathcal{F}(\hat{\theta})}$ (25)
@@
If the state of interest is pure, the quantum Fisher information is easily given by $4\langle\Delta\hat{H}\rangle$, which is proportional to $4\langle\Delta\hat{n}\rangle$ for optical systems, where $\hat{n}$ is the number operator $\hat{a}^{\dagger}\hat{a}$. For simplicity, we will call $4\langle\Delta\hat{n}\rangle$ as quantum Fisher information from now on. To begin with, quantum Fisher information of HESs is
$\mathcal{F}_{hyb}=4\alpha^{2},$ (26)
@@
which is again the same as the coherent state due to Proposition 1. In the same argument of Ref. Park16 , this supports that quantum Fisher information is a good measure of the overall performance of amplification schemes because both amplitude gain and noiseless property imply larger quantum Fisher information. Also, quantum Fisher information of $\hat{a}\hat{a}^{\dagger}$-amplified hybrid states is given by
$\mathcal{F}_{hyb}^{\hat{a}\hat{a}^{\dagger}}=\frac{4\alpha^{2}(\alpha^{8}+6\alpha^{6}+14\alpha^{4}+10\alpha^{2}+4)}{(\alpha^{4}+3\alpha^{2}+1)^{2}}.$ (27)
lastly, quantum Fisher information of $(\hat{a}^{\dagger})^{2}$-amplified hybrid states is given by
$\mathcal{F}_{hyb}^{\hat{a}^{\dagger}{}^{2}}=\frac{4\alpha^{2}(\alpha^{8}+8\alpha^{6}+24\alpha^{4}+24\alpha^{2}+12)}{(\alpha^{4}+4\alpha^{2}+2)^{2}}.$ (28)
@@
The ratios of quantum Fisher information, $\mathcal{F}_{hyb}^{\hat{a}\hat{a}^{\dagger}}/\mathcal{F}_{hyb}$ and $\mathcal{F}_{hyb}^{\hat{a}^{\dagger}{}^{2}}/\mathcal{F}_{hyb}$, are plotted in Fig 1. (c). The value is always larger than $1$ which indicates that hybrid states never become worse in terms of quantum Fischer information. In the asymptotic limit, the ratio converges to $1$, implying that the amplification scheme is of no use for sufficiently large amplitude.
@@
For a clear comparison of both schemes, we examined the ratio between two schemes $\mathcal{F}_{hyb}^{\hat{a}\hat{a}^{\dagger}}/\mathcal{F}_{hyb}^{\hat{a}^{\dagger}{}^{2}}$, in Fig 1. (d). The figure shows that up to amplitude $\alpha\approx 0.9$, $\hat{a}\hat{a}^{\dagger}$ scheme has higher quantum Fisher information. The ratio monotonically decreases up to $\alpha\approx 1.43$, starting from $1.33$ at $\alpha=0$, $1.11$ at $\alpha=0.45$, and $1$ at $\alpha\approx 0.9$. After $\alpha\approx 0.9$, $(\hat{a}^{\dagger})^{2}$ scheme yields higher quantum Fisher information for the amplified state. For sufficiently large $\alpha$, the ratio converges to $1$ and the difference is negligible. This result confirms our qualitative analysis.
@@ IV Amplification of Superpositions of Coherent States
In this section, we compare the two amplification schemes performed on SCSs. After $\hat{a}\hat{a}^{\dagger}$ scheme and $(\hat{a}^{\dagger})^{2}$ scheme, SCSs are amplified to
$N_{SCS}^{\hat{a}\hat{a}^{\dagger}}(\alpha,k,d)\,\hat{a}{\hat{a}^{\dagger}}\sum_{n}\omega_{d}^{-kn}\ket{\alpha\omega_{d}^{n}},$ (29)
and
$N_{SCS}^{\hat{a}^{\dagger}{}^{2}}(\alpha,k,d)\,(\hat{a}^{\dagger})^{2}\sum_{n}\omega_{d}^{-kn}\ket{\alpha\omega_{d}^{n}},$ (30)
respectively, where $N_{SCS}^{\hat{a}\hat{a}^{\dagger}}(\alpha,k,d)$ and $N_{SCS}^{\hat{a}^{\dagger}{}^{2}}(\alpha,k,d)$ are normalization factors
@@
$N_{SCS}^{\hat{a}^{\dagger}{}^{2}}(\alpha,k,d)=\frac{1}{\sqrt{d\sum_{n}\omega_{d}^{-kn}(\alpha^{4}\omega_{d}^{2n}+4\alpha^{2}\omega_{d}^{n}+2)\mathrm{Exp}[-\alpha^{2}(1-\omega_{d}^{n})]}}.$ (32)
One can see that the normalization factors $N_{SCS}^{\hat{a}^{\dagger}{}^{2}}(\alpha,k,d)$ and $N_{SCS}^{\hat{a}^{\dagger}{}^{2}}(\alpha,k,d)$ both depend on the qudit number $k$ and dimension $d$, unlike hybrid states.
In the same spirit as Sec. III, we first obtain fidelity and then obtain gain which is defined by the value that maximizes fidelity. Quantum Fisher information will be employed as an example of the overall performance of the two schemes. For each scheme, fidelity is written as
@@
Gain values are given by which maximize the fidelity, by solving following equations.
@@
This cannot be obtained using elementary functions in closed forms, thus we find the numerical solution. Relations between fidelity and gain for both schemes are shown in Fig 2. and Fig 3. for $d=3,4,5$ with every $k$. In Fig 2., fidelity and gain of the $\hat{a}\hat{a}^{\dagger}$ amplification are plotted with respect to the amplitude. Overall shapes of the graphs are not much different from those of hybrid states, however, there are two additional features. (i) The larger the qudit number $k$, the larger the amplitude $\alpha$ where the local minimum appears. (ii) The larger the qudit number $k$, the larger the local minimum.
Figure 2: For SCSs, $\hat{a}\hat{a}^{\dagger}$ scheme, (a) Fidelity plot of d=3. (b) Optimal gain plot of d=3. (c) Fidelity plot of d=4. (d) Optimal plot of d=4. (e) Fidelity plot of d=5. (f) Optimal gain plot of d=5.
Recall that SCSs with qudit number $k$, |$\ket{\mathcal{C}^{k}_{\alpha,d}}$, contains $k(\mathrm{mod}\ d)$ number of photons. For small amplitude, $\alpha\approx 0$, the coefficient of Fock basis |$\ket{n}$ is concentrated on the Fock state with the least number, |$\ket{k}$. This explains high fidelity for very low amplitude because the amplification $\hat{a}\hat{a}^{\dagger}$ does not change the number state, as one can confirm with the following equality
@@
As $\alpha$ increases, the SCSs no longer stay in the number state $k$. |$\ket{\mathcal{C}^{k}_{\alpha,d}}$ contains more and more |$\ket{d+k}$, |$\ket{2d+k}$, …, Fock states. Then, the additional factor arising from the amplification becomes relevant and changes the state.
@@
We revisit the pseudo-number property of SCSs to further investigate the behavior of their fidelities against amplified counterparts. First observe from Eq. (9) that as the amplitude $\alpha$ increases, the mean photon number of SCSs grows. While the physical interpretation of the amplitude $\alpha$ renders this intuitive, one can verify it by examining the sign of the derivative. As we mentioned earlier, |$\ket{\mathcal{C}^{k}_{\alpha,d}}$ starts from the Fock state |$\ket{k}$ for $\alpha\sim 0$, and obtain coefficients of other Fock states |$\ket{d+k}$, |$\ket{2d+k}$, $\dots$, as $\alpha$ increases. For small qudit-number $k$, coefficients of other Fock states can be sufficiently large to reduce the fidelity with small amplitude $\alpha$. This explains why trough appears earlier with smaller qudit-number $k$. As amplitude $\alpha$ grows, overlap between {|$\{\ket{\alpha\omega_{d}^{n}}\}_{n}$ becomes smaller, thus recovered orthogonality contributes to fidelity which makes the depth of trough with larger $k$ shallower than that of smaller $k$.
@@
With this qualitative analysis, we can predict two additional features that were not seen in Fig 2. First, since the overlap between {|$\{\ket{\alpha\omega_{d}^{n}}\}_{n}$ increases with $d$, we need larger amplitude $\alpha$ to recover orthogonality. Then the coefficients of |$\ket{2d+k},\ket{3d+k},\dots$, can be significantly large before we recover orthogonality, thus it makes multiple troughs, as we showed in Fig 3. Second, for fixed $k$, there is no monotonic tendency of the depth of troughs concerning $d$ because two factors antagonize each other. If $d$ grows, overlap between {|$\{\ket{\alpha\omega_{d}^{n}}\}_{n}$ increases, but the required amplitude $\alpha$ to see trough becomes large, contributing to orthogonality. One can see that the depth of troughs of qudit number 00 with $d=3$ is deeper than that of $d=2$ and $d=4$.
@@
The gain graphs do not show any particular behavior. As $\alpha$ increases, more and more photons SCSs have, thus contribution to the gain of $\hat{a}\hat{a}^{\dagger}$ scheme monotonically decreases with $\alpha$.
Figure 3: For SCSs, $\hat{a}^{\dagger}{}^{2}$ scheme, (a) Fidelity graphs of d=3. (b) Gain graphs of d=3. (c) Fidelity graphs of d=4. (d) Gain graphs of d=4. (e) Fidelity graphs of d=5. (f) Gain graphs of d=5.

### v2

@@ III Amplification of Hybrid Entangled states
For the $\hat{a}\hat{a}^{\dagger}$ scheme, solving the equation gives the optimized gain as
$G_{HES}^{\hat{a}\hat{a}^{\dagger}}=\frac{\alpha^{2}-1+\sqrt{\alpha^{4}+6\alpha^{2}+1}}{2\alpha^{2}},$ (21)
while for the $(\hat{a}^{\dagger})^{2}$ scheme, the optimized gain is obtained as
$G_{HES}^{\hat{a}^{\dagger}{}^{2}}=\frac{1+\sqrt{1+\frac{8}{\alpha^{2}}}}{2}.$ (22)
The optimized gains of the two schemes are shown in Fig. 1(b). One can see that the $\hat{a}^{\dagger}{}^{2}$ scheme yields the higher optical gain than the $\hat{a}\hat{a}^{\dagger}$ scheme in contrast with the fidelities. The behavior of optimized gain of hybrid state amplification is also identical to that of coherent state amplification for the same reason as the case of the fidelity.
We further note that the gain can be defined by the ratio of the expectation values of quadrature operators with an arbitrary phase $\phi$ Zavatta11 ; Park16 . This gain value, however, can be somewhat inadequate for HESs because the expectation value of quadrature value with any phase $\phi$ is zero for HESs, as one can confirm with the equation
@@
$g=(N^{f(\hat{a},{\hat{a}^{\dagger}})})^{2}\frac{\bra{\alpha}f(\hat{a},{\hat{a}^{\dagger}})^{\dagger}(\hat{a}e^{i\lambda}+\hat{a}^{\dagger}e^{-i\lambda})f(\hat{a},{\hat{a}^{\dagger}})\ket{\alpha}}{\bra{\alpha}(\hat{a}e^{i\lambda}+\hat{a}^{\dagger}e^{-i\lambda})\ket{\alpha}}.$
If we write the normal ordered form of $f(\hat{a},{\hat{a}^{\dagger}})^{\dagger}\hat{a}f(\hat{a},{\hat{a}^{\dagger}})$ as $\sum_{i}C_{i}\,\hat{a}^{\dagger}{}^{i}\hat{a}^{\dagger}{}^{i}\hat{a}=g(\hat{a},\hat{a}^{\dagger})\hat{a}$, the gain is expressed as
$g=(N^{f(\hat{a},{\hat{a}^{\dagger}})})^{2}\bra{\alpha}g(\hat{a},\hat{a}^{\dagger})\ket{\alpha}.$
Likewise, the gain of HESs defined by the ratio is
$g_{HES}=(N_{HES}^{f(\hat{a},{\hat{a}^{\dagger}})})^{2}\bra{\alpha}g(\hat{a},\hat{a}^{\dagger})\ket{\alpha}.$
@@
We can go further with this gain to define the EIN Grangier92 , a measure of noiseless amplification. However, since the expectation value of quadrature is zero, it lacks physical meaning, and we will not stick to it.
@@
Observing the results so far, the $\hat{a}\hat{a}^{\dagger}$ scheme has an advantage of higher fidelity while $\hat{a}^{\dagger}{}^{2}$ has an advantage of higher gain, for any amplitude $\alpha$. This property provides us with a guide of which amplification scheme should be used. However, their advantages in different aspects make it hard to compare their overall performance. Qualitatively, we know that the $\hat{a}\hat{a}^{\dagger}$ scheme would be better for small amplitudes since coherent states with small amplitudes are susceptible to the change of their photon numbers. Meanwhile, we can guess that the two schemes will not differ so much for large amplitudes because coherent states with sufficiently large amplitudes are robust against several photon addition or subtraction processes. Lastly, as a consequence, we can deduce that the $\hat{a}^{\dagger}{}^{2}$ scheme would be appropriate for the intermediate amplitude region.
@@
A previous work Park16 employed two measures, the EIN for coherent states and the quantum Fisher information for $2$-dimensional SCSs to quantify overall performance which reflects aforementioned qualitative arguments. The EIN is a measure of undesired noise generated in the amplification process, thus it is appropriate to measure overall performance. However, we have concluded that the EIN lacks physical meaning in our case since the expectation values of quadrature operators are always zero.
As an alternative, we take the quantum Fisher information associated with the Hamiltonian of the harmonic oscillator $\hat{H}$. The quantum Fisher information is then defined as
$\mathcal{F}[\hat{\rho},\hat{H}]=2\sum_{k,l}\frac{(\lambda_{k}-\lambda_{l})^{2}}{(\lambda_{k}+\lambda_{l})}|\bra{k}\hat{H}\ket{l}|^{2},$ (24)
@@
where $\hat{H}:=\hat{a}^{\dagger}\hat{a}$, $\hat{\rho}$ is a density matrix of a state used for the estimation of parameter $\theta$, |$\ket{k}$ is a $k$-th eigenvector, and $\lambda_{k}$ is a $k$-th eigenvalue of the density matrix $\rho$. Here, we use a parametrized density matrix $\hat{\rho}(\theta):=e^{-i\hat{H}\theta}\hat{\rho}e^{i\hat{H}\theta}$ for the phase estimation. We only consider the bosonic mode even for the HES to make a fair comparison. The quantum Fisher information is in a direct relationship with the performance of phase estimation. The quantum Cramér-Rao bound,
$\mathrm{var}(\theta)\geq\frac{1}{\mathcal{F}(\theta)},$ (25)
imposes a lower bound on the statistical error of the phase estimation scheme.
@@
If the state of interest is pure, the quantum Fisher information is easily given by $4\langle\Delta\hat{H}\rangle$, which is proportional to $4\langle\Delta\hat{n}\rangle$ for optical systems, where $\hat{n}$ is the number operator $\hat{a}^{\dagger}\hat{a}$. For simplicity, we will call $4\langle\Delta\hat{n}\rangle$ as the quantum Fisher information from now on. To begin with, the quantum Fisher information of the HES is
$\mathcal{F}_{HES}=4\alpha^{2},$ (26)
@@
which is again the same as that of the coherent state due to Proposition 1. In the same argument of Ref. Park16 , this supports that the quantum Fisher information is a good measure of the overall performance of the amplification schemes because both amplitude gain and noiseless property imply larger quantum Fisher information. The quantum Fisher information of $\hat{a}\hat{a}^{\dagger}$-amplified hybrid states is also given by
$\mathcal{F}_{HES}^{\hat{a}\hat{a}^{\dagger}}=\frac{4\alpha^{2}(\alpha^{8}+6\alpha^{6}+14\alpha^{4}+10\alpha^{2}+4)}{(\alpha^{4}+3\alpha^{2}+1)^{2}},$ (27)
and the quantum Fisher information of $\hat{a}^{\dagger}{}^{2}$-amplified hybrid states is
$\mathcal{F}_{HES}^{\hat{a}^{\dagger}{}^{2}}=\frac{4\alpha^{2}(\alpha^{8}+8\alpha^{6}+24\alpha^{4}+24\alpha^{2}+12)}{(\alpha^{4}+4\alpha^{2}+2)^{2}}.$ (28)
@@
The ratios of quantum Fisher information, $\mathcal{F}_{HES}^{\hat{a}\hat{a}^{\dagger}}/\mathcal{F}_{HES}$ and $\mathcal{F}_{HES}^{\hat{a}^{\dagger}{}^{2}}/\mathcal{F}_{HES}$, are plotted in Fig. 1(c). The value is always larger than $1$, indicating that the HESs never become worse in terms of the quantum Fisher information by the amplification schemes. In the asymptotic limit, the ratio converges to $1$, implying that the amplification schemes are trivial in such a regime.
@@
For a clear comparison of both schemes, we examined the ratio between two schemes $\mathcal{F}_{HES}^{\hat{a}\hat{a}^{\dagger}}/\mathcal{F}_{HES}^{\hat{a}^{\dagger}{}^{2}}$, in Fig. 1(d). The figure shows that up to amplitude $\alpha\approx 0.9$, the $\hat{a}\hat{a}^{\dagger}$ scheme has the higher quantum Fisher information. The ratio monotonically decreases to $\alpha\approx 1.43$, starting from $1.33$ at $\alpha=0$, $1.11$ at $\alpha=0.45$, and $1$ at $\alpha\approx 0.9$. After $\alpha\approx 0.9$, the $\hat{a}^{\dagger}{}^{2}$ scheme yields the higher quantum Fisher information for the amplified state. For sufficiently large $\alpha$, the ratio converges to $1$ and the difference is negligible. This result confirms our qualitative analysis.
@@ IV Amplification of Superpositions of Coherent States
In this section, we compare the two amplification schemes performed on SCSs. After the $\hat{a}\hat{a}^{\dagger}$ scheme and $\hat{a}^{\dagger}{}^{2}$ scheme, respectively, the amplified SCSs can be expressed as
|$\ket{\mathcal{C}_{\alpha,d}^{k,\hat{a}\hat{a}^{\dagger}}}:=N_{SCS}^{\hat{a}\hat{a}^{\dagger}}(\alpha,k,d)\,\hat{a}{\hat{a}^{\dagger}}\sum_{n}\omega_{d}^{-kn}\ket{\alpha\omega_{d}^{n}},$ (29)
|$\ket{\mathcal{C}_{\alpha,d}^{k,(\hat{a}^{\dagger})^{2}}}:=N_{SCS}^{(\hat{a}^{\dagger}{})^{2}}(\alpha,k,d)\,(\hat{a}^{\dagger})^{2}\sum_{n}\omega_{d}^{-kn}\ket{\alpha\omega_{d}^{n}},$
where $N_{SCS}^{\hat{a}\hat{a}^{\dagger}}(\alpha,k,d)$ and $N_{SCS}^{(\hat{a}^{\dagger}{})^{2}}(\alpha,k,d)$ are normalization factors
@@
$N_{SCS}^{(\hat{a}^{\dagger}{})^{2}}(\alpha,k,d)=\frac{1}{\sqrt{d\sum_{n}\omega_{d}^{-kn}(\alpha^{4}\omega_{d}^{2n}+4\alpha^{2}\omega_{d}^{n}+2)\mathrm{exp}[-\alpha^{2}(1-\omega_{d}^{n})]}}.$ (31)
Both the normalization factors $N_{SCS}^{(\hat{a}^{\dagger}{})^{2}}(\alpha,k,d)$ and $N_{SCS}^{(\hat{a}^{\dagger}{})^{2}}(\alpha,k,d)$ depend on the qudit number $k$ and dimension $d$, unlike the normalization factors of the amplified HESs. Explicit forms of the amplified SCSs are presented in Appendix B.
As explained in Sec. III with Eq. (15), we first obtain the fidelities and the optimized gains defined by the values that maximize the fidelities:
$F^{\hat{a}\hat{a}^{\dagger}}_{SCS}:=\absolutevalue{\innerproduct{\mathcal{C}^{k}_{g\alpha,d}}{\mathcal{C}^{k,\hat{a}\hat{a}^{\dagger}}_{\alpha,d}}}^{2},$ (32)
$F^{\hat{a}^{\dagger}{}^{2}}_{SCS}:=\absolutevalue{\innerproduct{\mathcal{C}^{k+2}_{g\alpha,d}}{\mathcal{C}^{k,(\hat{a}^{\dagger})^{2}}_{\alpha,d}}}^{2}.$
For each scheme, the fidelity can be expressed as a function of $\alpha$ and $g$ as
@@
The optimized gains are then obtained to maximize the fidelities:
@@
We find numerical values of the gain using the downhill simplex algorithm since analytic forms cannot be obtained in closed forms using elementary functions. The implementation was performed through the Python package Scipy, using the  scipy fmin function. Relations between the fidelity and gain for both the schemes are shown in Figs. 2 and 3 for $d=3,4,5$ with every $k$. In Fig. 2, the fidelity and the optimized gain of the $\hat{a}\hat{a}^{\dagger}$ amplification on SCSs are plotted against the coherent amplitude $\alpha$. Behaviors of those quantities do not differ much from the HES case, however, there are two additional features. (i) The larger the qudit number $k$, the larger the amplitude $\alpha$ where the local minimum appears. (ii) The larger the qudit number $k$, the larger the local minimum.
Figure 2: Fidelity (F) and optimized gain (G) obtained from the $\hat{a}\hat{a}^{\dagger}$ scheme against input coherent amplitude $\alpha$ of SCSs with qudit dimension [(a), (b) $d=3$, (c),(d) $d=4$, (e), (f) $d=5$], respectively. Here, $k$ is the integer qudit index in Eq. 7 with the relative phase information of the $d$-dimensional qudit.
Figure 3: Fidelity (F) and optimized gain (G) obtained from the $\hat{a}^{\dagger}{}^{2}$ scheme against coherent amplitude $\alpha$ of input SCSs with qudit dimension [(a), (b) $d=3$, (c), (d) $d=4$, (e), (f) $d=5$], respectively.
We recall that a SCS with qudit number $k$, |$\ket{\mathcal{C}^{k}_{\alpha,d}}$, contains $k\leavevmode\nobreak\ (\mathrm{mod}\ d)$ number of photons. For a small amplitude, the coefficient of Fock basis |$\ket{n}$ is concentrated on the Fock state with the least number, |$\ket{k}$. This explains high fidelity for very low amplitude because the amplification $\hat{a}\hat{a}^{\dagger}$ does not change the number state, as one can confirm with the following equality
@@
As $\alpha$ increases, the SCS is no longer an approximation of a number state of $k$. State |$\ket{\mathcal{C}^{k}_{\alpha,d}}$ contains more Fock states such as |$\ket{d+k}$, |$\ket{2d+k}$, …. The additional factor arising from the amplification then becomes relevant and changes the state as
@@
We revisit the pseudo-number property of SCSs to further investigate the behavior of their fidelities against amplified counterparts. It is clear that the mean photon numbers of the SCSs become larger as the amplitudes $\alpha$ increase as also seen in Eq. (9). While the physical interpretation of the amplitude $\alpha$ renders this intuitive, one can verify it by examining the sign of the derivative. As we mentioned earlier, |$\ket{\mathcal{C}^{k}_{\alpha,d}}$ starts from the Fock state |$\ket{k}$ for $\alpha\sim 0$, and obtain coefficients of other Fock states |$\ket{d+k}$, |$\ket{2d+k}$, $\dots$, as $\alpha$ increases. For small qudit-number $k$, coefficients of other Fock states can be sufficiently large to reduce the fidelity with small amplitude $\alpha$. This explains why trough appears earlier with smaller qudit-number $k$. As amplitude $\alpha$ grows, the overlap between {|$\{\ket{\alpha\omega_{d}^{n}}\}_{n}$ becomes smaller, and the recovered orthogonality contributes to the fidelity which makes the depth of trough shallower as $k$ increases.
@@
With this qualitative analysis, we can predict two additional features that were not seen in Fig. 2. First, since the overlap between {|$\{\ket{\alpha\omega_{d}^{n}}\}_{n}$ increases with $d$, we need larger amplitude $\alpha$ to recover the orthogonality. The coefficients of |$\ket{2d+k},\ket{3d+k},\dots$, can then be significantly large before we recover the orthogonality, and it makes multiple troughs as shown in the fidelity plots of Fig. 3. Second, for fixed $k$, there is no monotonic tendency of the depth of troughs concerning $d$ because two factors antagonize each other. If $d$ grows, the overlap between {|$\{\ket{\alpha\omega_{d}^{n}}\}_{n}$ increases, but the required amplitude $\alpha$ to see trough becomes large, contributing to orthogonality. One can see that the depth of troughs of qudit number 00 with $d=3$ is deeper than that of $d=2$ and $d=4$.
@@
The gain graphs of Fig. 3 do not show any unusual behaviors. The SCSs have more photons as $\alpha$ increases, and the contribution to the gain for the $\hat{a}\hat{a}^{\dagger}$ scheme monotonicall decreases with $\alpha$.